There are an equal number of boys and girls in a class. If there were 6 fewer boys and 6 more girls, there would be twice as many girls as boys. How many students are in the class?
Answer by ankor@dixie-net.com(12701)

There are an equal number of boys and girls in a class. If there were 6 fewer boys and 6 more girls, there would be twice as many girls as boys. How many students are in the class?
:
Let b = no. of boys
&
g = no. of girls
:
(g + 6) = 2(b - 6)
g + 6 = 2b - 12
g = 2b - 12 - 6
g = 2b - 18
we know b = g, therefore:
b = 2b - 18
b - 2b = -18
-b = -18
b = 18 boys and 18 girls, a total of 36 in the class
:
:
check: (18+6) = 2(18-6)